Find the exact value of each of the remaining trigonometric functions of θ.
Tanθ=-root(35), cosθ<0.
sinθ=?
cosθ=?
secθ=?
cscθ=?
cotθ=?

Answer by lwsshak3(11628)

Find the exact value of each of the remaining trigonometric functions of θ.
Tanθ=-root(35), cosθ<0.
sinθ=?
cosθ=?
secθ=?
cscθ=?
cotθ=?
**
tanθ=-√35/1=opp side/adj side (in quadrant II where tan<0 and cos<0)
hypotenuse:
=√[(√35)^2+(1^2)]
=√(35+1)
=√36
=6
..
sinθ=√35/6
cosθ=-1/6
secθ=-6
cscθ=6/√35
cotθ=-1/√35